Mmabatho the Social Scientist took a random sample of 30 adults with Autism Spectrum Disorder (ASD) and found their reading time to be normally distributed with sample mean and sample
standard deviation of 90 wpm and 18 wpm respectively.

Question 5
What is the 99% confidence interval estimate of the population mean?
(1) .80:9626I 99:0374/:
(2) .80:9429I 99:0571/:
(3) .81:9090I 98:0910/:
(4) .81:5213I 98:4787/:
(5) .82:3428I 97:6572/:
1
Expert's answer
2020-03-27T16:42:15-0400

"99\%CI=(\bar x-t_{0.005,n-1}\frac{s}{\sqrt{n}},\bar x+t_{0.005,n-1}\frac{s}{\sqrt{n}})="

"=(90-2.7564\frac{18}{\sqrt{30}}, 90+2.7564\frac{18}{\sqrt{30}})="

"=(80.9429,99.0571)".

Var. (2)